OpenStudy (diana):

how do you solve cos^2 ((1/2)sin^1(3/5))?

OpenStudy (anonymous):

Do you mean \[\cos^2\left(\frac{1}{2}\sin^{-1}(\frac{3}{5})\right) \] ? Also, what do you mean by solve?

OpenStudy (diana):

yeah

OpenStudy (diana):

like solve it without a calculator

OpenStudy (anonymous):

Is there any particular reason that you can't use a calculator?

OpenStudy (anonymous):

basically, because calculator say "syntax error". . .

OpenStudy (anonymous):

1/10, is that the answer?